OpenStudy (anonymous):

which of the following is a solution to the equation y = 3x - 1 a. (4,1) b. (2,5) c. (4,3) d. (0,-3)

OpenStudy (ineedhelpfast12):

I believe the answer is A

jimthompson5910 (jim_thompson5910):

let's test that y = 3x - 1 1 = 3*4 - 1 ... plug in (x,y) = (4,1) 1 = 12 - 1 1 = 11 ... this is false, so (4,1) is NOT a solution to y = 3x - 1

OpenStudy (ineedhelpfast12):

hmmm I graphed it out and used 3x as the slope and -1 as the y-intercept, meaning that I put a dot on (0,-1) and then went up three and over 1, so the answer would be (2,-1). My bad!